I would like to sell 2 gallon cans of peaches. What would the diameter of the base of the can need to be for it to be 8 inches tall and able to hold 2 gallon(s) of liquid? (Note: 1 Gallon = 231 Cubic Inches)
Answer by Cromlix(4381)

Formula for volume of cylinder=
Pi*r^2*h
Vol = Pi*r^2 *h
462 = Pi*r^2*8
r^2 = 462/Pi*8
r^2 = 18.38
r = square root of 18.38
r = 4.29 inches
diameter = 8.57 inches.
Hope this helps.
:-)